a bag contains 4 red marbles, 5 blue marbles, and 3 green marbles. a randomly drawn marble is blue.
Answer #1

Bag contains Gred, 5 blue, 3 preen. Find the probability a randomly drawin marble is bleu. Total mambel = 4+ 5+ 32 12 ) looso
Answer #2

SOLUTION :


Total marbles in the bag = 4 + 5 + 3 = 12


P( blue) 

= blue marbles in the bag  / Total marbles in the bag

= 5/12  (ANSWER)

answered by: Tulsiram Garg